# An interpretable generative multimodal neuroimaging-genomics framework for decoding Alzheimer’s disease

### v1

@@ Abstract
Alzheimer’s disease (AD) is the most prevalent form of dementia, affecting millions worldwide with a progressive decline in cognitive abilities. The AD continuum encompasses a prodormal stage known as Mild Cognitive Impairment (MCI), where patients may either progress to AD (MCIc) or remain stable (MCInc).
Understanding the underlying mechanisms of AD requires complementary analysis derived from different data sources, leading to the development of multimodal deep learning models. In this study, we leveraged structural and functional Magnetic Resonance Imaging (sMRI/fMRI) to investigate the disease-induced grey matter and functional network connectivity changes. Moreover, considering AD’s strong genetic component, we introduce Single Nucleotide Polymorphisms (SNPs) as a third channel. Given such diverse inputs, missing one or more modalities is a typical concern of multimodal methods. We hence propose a novel deep learning based classification framework where generative module employing Cycle Generative Adversarial Networks (cGAN) was adopted to impute missing data within the latent space. Additionally, we adopted an Explainable Artificial Intelligence (XAI) method, Integrated Gradients (IG), to extract input features relevance, enhancing our understanding of the learned representations.
Two critical tasks were addressed: AD detection and MCI conversion prediction. Experimental results showed that our framework was able to reach the state-of-the-art in the classification of CN vs AD reaching an average test accuracy of $0.926\pm 0.02$. For the MCInc vs MCIc task, we achieved an average prediction accuracy of $0.711\pm 0.01$ using the pre-trained model for CN and AD. The interpretability analysis revealed that the classification performance was led by significant grey matter modulations in cortical and subcortical brain areas well known for their association with AD. Moreover, impairments in sensory-motor and visual resting state network connectivity along the disease continuum, as well as mutations in SNPs defining biological processes linked to amyloid-beta and cholesterol formation clearance and regulation, were identified as contributors to the achieved performance. Overall, our integrative deep learning approach shows promise for AD detection and MCI prediction, while shading light on important biological insights.
Keywords Alzheimer’s disease continuum  $\cdot$ Generative model  $\cdot$ Imaging genetics  $\cdot$ Integrated Gradients
@@ 1 Introduction
Among the available neuroimaging technologies, structural Magnetic Resonance Imaging (sMRI) and resting-state functional MRI (rs-fMRI) have provided unprecedented opportunities for deriving biomarkers allowing the early diagnosis of AD. For instance, sMRI is currently a key part of the diagnostic criteria for the differential diagnosis and longitudinal monitoring of patients with dementia, enabling the estimation of brain atrophy. Several studies have consistently observed both global and local atrophic changes in AD, lying along the hippocampal pathway (entorhinal cortex, hippocampus, parahippocampal gyrus, and posterior cingulate cortex) in the early stages of the disease, while atrophy in temporal, parietal and frontal neocortices emerges at later stages being associated with neuronal loss as well as with language, visuospatial and behavioral impairments [7, 8]. Rs-fMRI, in turn, indirectly measures the neural activity by detecting changes in the Blood Oxygenation Level Dependent (BOLD) signals, which depend on the neurovascular coupling [9]. In particular, investigating functional connectivity (FC; inter-regional coupling), functional network connectivity (FNC; inter-network coupling), and functional networks from BOLD rs-fMRI provides a means for understanding the mechanisms and relevance of the functional relationships across brain regions. A growing body of rs-fMRI studies suggests that failure of specific resting-state networks (RSNs) is closely related to AD, with the default-mode network (DM) and the salience network (SN) playing a pivotal role and being disrupted before clinically evident symptoms [10]. Specific alterations in these functional networks have been reported in AD patients, with prominent FC decreased within the DM and increased FC in the SN [11, 12]. Moreover, disconnections within and between the different RSNs have been consistently demonstrated, particularly over long connection distances [13]. All of these factors have contributed to the widespread view of AD as a disconnection syndrome being characterized by a cascading network failure, beginning in the posterior DM and then shifting to other systems containing prominent connectivity hubs, possibly associated with amyloid accumulation [14]. Moreover, increased evidence supports the view that tau depositions are also related to functional brain architecture and FC changes, supporting the view of transneuronal tau propagation in AD [15].
@@
Moreover, AD also features a strong genetic component. Strategies to extract linked genotype traits are commonly based on Single Nucleotide Polymorphism (SNPs) analysis [16], representing variations of single nucleotides in DNA sequences that vary from person to person [16] and are present in at least $1\%$ of the population. Several Genome-Wide Association Studies (GWAS) have identified more than 40 AD-associated genes/loci, which are likely to increase the risk of developing the disease [17, 18, 19]. Among them the apolipoprotein E (APOE) gene, in particular the $\epsilon 4$ allele, PICALM, CLU, ABCA7, and CR1111 are the most important genes being associated with AD risk factor or the progression from MCI to AD [20], [21], [22], [23], [24], [25], [26, 27].
@@
This great heterogeneity of available biomarkers offers a unique opportunity to explore various aspects of the disease continuum. Each biomarker provides valuable information about specific characteristics, enabling a multifaceted investigation of AD which calls for methods that can effectively integrate and leverage complementary information. In this regard, artificial intelligence, particularly Deep Learning (DL), emerges as a promising technology to tackle this complex task. By employing multiple layers of processing, DL models allow extracting progressively high-level and more informative features from input data. Moreover, the inclusion of multiple data sources into a single model can uncover complex and deep non-linear associations between the input features from a multimodal perspective [28]. As a result, multimodality is gaining significant popularity representing the key approach to gain valuable insights into complex and multifaceted neurodegenerative diseases such as AD [29], [30], [31]. However, despite the promising foreseen of such an approach, multiple drawbacks are present and represent the focus of the current research. The main concern resides in missing data management. Particularly in the biomedical domain, it is very common to incur in missing values or acquisitions for certain subjects or entire study cohorts due to different reasons, such as missing acquisition, corrupted data, and patients dropout from a study [32], privacy and expensive tests. Additionally, the interpretability of a model’s predictions is a central characteristic of decision-aiding models, which is still not pervasively addressed. In the past years, many high-performing prediction models have been proposed lacking a clear rationale to be effectively considered for practical use. Strong and validated explanations associated with a given prediction are fundamental for increasing trust in the results as well as their applicability in a real-world scenario. Innovative and viable missing modality management as well as interpretability are the key attributes that a multimodal model for diagnosis detection should have.
@@
Diving into missing data management, the simplest and most common solutions consist either of discarding samples with missing modalities, or filling the missing values with zeros [33], and computing imputations based on data interpolation [34]. Such solutions are evidently suboptimal since they could significantly reduce the number of training samples, already small when addressing biomedical-related tasks, or introduce important biases in the data and the model due to the interpolation or the zero filling. Alternative methods to exploit the information of all the available subjects were proposed. A possible approach is the complementation of incomplete data representation, which consists in extracting a latent representation from both the complete and incomplete modalities, avoiding the need for data imputation [35, 36]. Most recent approaches aim at generating missing data either in the input space [37] or in an intermediate latent representation [38, 39] relying on generative models such as Generative Adversarial Networks (GANs) [40] and their variants, or Variational Autoencoders [41]. This last approach has been successfully applied to the AD continuum investigation. It allows the exploitation of the availability of multimodal data for capturing the relationship among different data sources in the latent space, enabling the generation of one modality from another. Generative models, hence, appear as the route to be followed when dealing with missing data. Interestingly, the recently proposed Cycle-consistent GANs (Cycle-GANs) [42] have shown impressive results in various knowledge translation tasks since they offer a flexible and effective approach for learning mappings across different domains without relying on paired data.
@@
Moving to model interpretability, it is well established that with increasing model complexity, interpretability decreases drastically. However, in order to better understand the mechanisms that underlie the AD continuum and allow the models to be applied in clinical and real-life applications, it is vital to understand the reasons behind a certain output. In this case, Explainable Artificial Intelligence (XAI), becomes essential to understand why a given model made a certain prediction. XAI encompasses post-hoc methods that allow the assignment of an importance score to each feature, reflecting its role in the classification task. This means finally opening the ‘black box’ of complex models hopefully increasing their exploitability in clinical practice. XAI is starting to be applied in multimodal frameworks to study neurodegenerative and psychiatric diseases. Few works could be found adopting simple gradient-based or feature perturbation methods [43, 44, 45]. On top of this, another overlooked yet stringent aspect is the strong validation of XAI methods and the obtained explanations. The evaluation of explanation methods is still under-investigated, however, since explainability is meant to increase confidence in AI, it is vital to systematically analyze the obtained results referring also to the prior knowledge derived from the state-of-the-art.
@@
In this rapidly evolving landscape, where multimodality holds a central role, we aim at shading light on the importance of the complementary information that could be derived from advanced biomarkers such as brain FC and SNPs mutations, together with the well-established brain atrophy measures derived from sMRI in detecting AD-related modulations. With this purpose, we will present a novel multimodal framework for AD detection and MCI conversion prediction which addresses (i) heterogeneous data integration (ii) missing data management, and (iii) interpretability. Regarding the former, our framework allows the integration of heterogeneous data featuring different dimensionality, e.g., 3D volumes and vectorized data, and nature, meaning from multi-domain, e.g., imaging and genetics. Concerning the second, we propose an approach that allows generating missing modalities in the latent space obtained after input feature reduction, ensuring high accuracy in reconstructing latent features while minimizing computational demands. The goal is to define a framework that can be generalized to any missing modality, without requiring to have at least one specific modality shared by all the subjects in the considered population. Finally, we aimed at emphasizing the interpretability of the proposed framework in order to reinforce its transparency and reliability by conducting a post-hoc interpretability analysis, supplemented by a robust validation step, which enables to precisely discern the contribution of each input feature to the classification of subjects in the AD continuum.
2 Material and Methods
@@ 2.1 Dataset
An accurate and extensive subject selection was performed to retain, for each subject, the baseline timepoint for all the selected imaging modalities, that is sMRI and rs-fMRI. The respective genetic variants (SNPs) were then selected, if available. The final study cohort included $1911$ subjects, divided into healthy controls (CN), AD, MCI non-converter (MCInc), and MCI converter (MCIc). In detail, following the classification available on the ADNI website, an MCI subject was considered as MCIc if dementia was diagnosed at any timepoint after MCI diagnosis.
Table 1 shows the AD and CN study cohorts demographic information highlighting the subjects sharing modalities, in particular, sMRI-fMRI and sMRI-SNPs. Table 2 details the MCI study cohort highlighting the percentage of missing data for each considered modality.
Table 1: Demographic information of the AD and CN study cohort. MMSE = Mini-Mental State Examination; wild (W) and mutated (M) alleles in APOE e4: 00 = homozygous (W/W), $1$ = heterozygous (W/M), $2$ = homozygous (M/M).
CN AD Cohorts All subjects sMRI and fMRI sMRI and SNPs All subjects sMRI and fMRI sMRI and SNPs Count $644$ $321$ $253$ $332$ $66$ $152$ Age (y) $73.6\pm 6.6$ $72.4\pm 7.1$ $76.4\pm 5.4$ $75.1\pm 7.9$ $74.7\pm 8.1$ $75.8\pm 7.7$ Gender (M/F) $284/360$ $117/204$ $137/116$ $181/151$ $35/31$ $84/68$ MMSE $29.1\pm 1.1$ $29.1\pm 1.2$ $29.1\pm 1.1$ $23.1\pm 2.2$ $22.6\pm 2.6$ $23.3\pm 2.0$ APOE e4444 ($0/1/2$) $455/168/21$ $224/87/10$ $189/57/7$ $109/156/67$ $19/34/13$ $49/76/27$
Table 2: Demographic information and missing data percentage of the MCI study cohort. MMSE = Mini-Mental State Examination; wild (W) and mutated (M) alleles in APOE4: 00 = homozygous (W/W), $1$ = heterozygous (W/M), $2$ = homozygous (M/M).
MCInc MCIc
Count 646 289
Age (y) $73\pm 8.5$ $74\pm 8.5$
Gender (M/F) $374$/$272$ $172$/$117$
MMSE $27.9\pm 1.8$ $26.9\pm 1.8$
APOE4 ($0/1/2$) $377/205/64$ $103/143/43$
Missing SNPs (%) $59.6\%$ $30.8\%$
Missing fMRI (%) $60.1\%$ $92.0\%$
3D T1-w MRI and rs-fMRI acquisitions were considered as imaging input channels and were acquired with the following sequence parameters: $1$) sagittal accelerated MPRAGE, TR/TE = shortest, TI = $900$ ms, flip angle = $9^{\circ}$, Field Of View = $256\times 256$ mm2, spatial resolution = $1\times 1\times 1$ mm3, slices = $176-211$), $2$) rs-fMRI: TR/TE =3000/303000303000/303000 / 30 ms, FA = $90^{\circ}$, FOV = $220\times 220\times 163$ mm3, $3.4$-mm isotropic voxel size. $200$ fMRI volumes were acquired in almost all subjects, with minimal variations in a small subset (e.g., $197$ or $195$ volumes). More details about the data acquisition can be found in [46]. Concerning the genetic data, DNA samples were genotyped using Illumina Human610-Quad or Illumina HumanOmniExpress BeadChip.

### v3

@@ Abstract
Objective: Alzheimer’s disease (AD) is the most prevalent form of dementia worldwide, encompassing a prodromal stage known as Mild Cognitive Impairment (MCI), where patients may either progress to AD or remain stable. The objective of the work was to capture structural and functional modulations of brain structure and function relying on multimodal MRI data and Single Nucleotide Polymorphisms, also in case of missing views, with the twofold goal of classifying AD patients versus healthy controls and detecting MCI converters. Approach: We propose a multimodal DL-based classification framework where a generative module employing Cycle Generative Adversarial Networks was introduced in the latent space for imputing missing data (a common issue of multimodal approaches). Explainable AI method was then used to extract input features’ relevance allowing for post-hoc validation and enhancing the interpretability of the learned representations. Main results: Experimental results on two tasks, AD detection and MCI conversion, showed that our framework reached competitive performance in the state-of-the-art with an accuracy of $0.926\pm 0.02$ and $0.711\pm 0.01$ in the two tasks, respectively. The interpretability analysis revealed gray matter modulations in cortical and subcortical brain areas typically associated with AD. Moreover, impairments in sensory-motor and visual resting state networks along the disease continuum, as well as genetic mutations defining biological processes linked to endocytosis, amyloid-beta, and cholesterol, were identified. Significance: Our integrative and interpretable DL approach shows promising performance for AD detection and MCI prediction while shedding light on important biological insights.
Keywords Alzheimer’s disease  $\cdot$ Generative model  $\cdot$ Imaging-genetics  $\cdot$ Explainable AI
@@ 1 Introduction
Among the available neuroimaging technologies, structural Magnetic Resonance Imaging (sMRI) and resting-state functional MRI (rs-fMRI) have provided unprecedented opportunities for deriving biomarkers allowing the early diagnosis of AD. For instance, sMRI is currently a key part of the diagnostic criteria for the differential diagnosis and longitudinal monitoring of patients with dementia, enabling the estimation of brain atrophy. Several studies have consistently observed both global and local atrophic changes in AD, lying along the hippocampal pathway (entorhinal cortex, hippocampus, parahippocampal gyrus, and posterior cingulate cortex) in the early stages of the disease, while atrophy in temporal, parietal and frontal neocortices emerges at later stages being associated with neuronal loss as well as with language, visuospatial and behavioral impairments [7, 8]. Rs-fMRI, in turn, indirectly measures neural activity by detecting changes in the Blood Oxygenation Level Dependent (BOLD) signals, which depend on the neurovascular coupling [9]. In particular, investigating functional connectivity (FC; inter-regional coupling), functional network connectivity (FNC; inter-network coupling), and functional networks from BOLD rs-fMRI provides a means for understanding the mechanisms and relevance of the functional relationships across brain regions. A growing body of rs-fMRI studies suggests that failure of specific resting-state networks (RSNs) is closely related to AD, with the default-mode network (DM) and the salience network (SN) playing a pivotal role and being disrupted before clinically evident symptoms [10]. Specific alterations in these functional networks have been reported in AD patients, with prominent FC decreased within the DM and increased FC in the SN [11, 12]. Moreover, disconnections within and between the different RSNs have been consistently demonstrated, particularly over long connection distances [13]. All of these factors have contributed to the widespread view of AD as a disconnection syndrome being characterized by a cascading network failure, beginning in the posterior DM and then shifting to other systems containing prominent connectivity hubs, possibly associated with amyloid accumulation [14]. Moreover, increased evidence supports the view that tau depositions are also related to functional brain architecture and FC changes, supporting the view of transneuronal tau propagation in AD [15].
@@
AD also features a strong genetic component. Strategies to extract linked genotype traits are commonly based on Single Nucleotide Polymorphism (SNPs) analysis [16], representing variations of single nucleotides in DNA sequences that vary from person to person [16] and are present in at least $1\%$ of the population. Several Genome-Wide Association Studies (GWAS) have identified more than 40 AD-associated genes/loci, which are likely to increase the risk of developing the disease [17, 18, 19]. Among them, the apolipoprotein E (APOE), in particular the $\epsilon 4$ allele, PICALM, CLU, ABCA7, and CR1111 are the most important genes being associated with AD risk factor or the progression from MCI to AD [20, 21, 22, 23, 24, 25, 26, 27].
@@
This great heterogeneity of available biomarkers offers a unique opportunity to explore various aspects of the disease continuum. Each biomarker provides valuable information about specific characteristics, enabling a multifaceted investigation of AD which calls for methods that can effectively integrate and leverage complementary information. In this regard, artificial intelligence, Machine Learning (ML) and particularly Deep Learning (DL), emerge as promising technologies to tackle this complex task. Different ML algorithms have been developed for integrating heterogeneous data, such as Canonical Correlation Analysis (CCA), Partial Least Squares (PLS), and ensemble methods, e.g. consisting of different Support Vector Machines (SVM). Even if these methods can perform well in studying neurodegenerative diseases [28, 29], they have some limitations, such as requiring vectorized input data, hence denaturing the spatial information of 3D volumes (e.g., sMRI or PET 3D images), and potentially being too simple for solving complex tasks. On the other hand, by employing multiple layers of processing, DL models allow extracting progressively higher-level and more informative features from input data, with the option of preserving spatial information, without the need of vectorizing the input features. Moreover, the inclusion of multiple data sources into a single model can uncover complex and deep non-linear associations across the input features from a multimodal perspective [30]. As a result, multimodality is gaining significant popularity representing the key approach to derive valuable insights into complex and multifaceted neurodegenerative diseases such as AD [31, 32, 33, 34, 35, 36]. However, despite the promising foreseen of such an approach, multiple drawbacks are present and represent the focus of the current research. The main concern resides in missing data management. Particularly in the biomedical domain, it is very common to incur in missing values or acquisitions for certain subjects or entire study cohorts due to different reasons, such as missing acquisition, corrupted data, patients dropout from a study [37], and the requirements of privacy and expensive tests. Additionally, the interpretability of a model’s predictions is a central characteristic of decision-aiding models, which is still not pervasively addressed. In the past years, many high-performing prediction models have been proposed lacking a clear rationale to be effectively considered for practical use. Strong and validated explanations associated with a given prediction are fundamental for increasing trust in the results as well as for their applicability in a real-world scenario. Innovative and viable missing modality management as well as interpretability are the key attributes that a multimodal model for diagnosis detection should have.
@@
Diving into missing data management, the simplest and most common solutions consist either in discarding samples with missing modalities, or in filling the missing values with zeros [38], and computing imputations based on data interpolation [39]. Such solutions are evidently suboptimal since they could significantly reduce the number of training samples, already small when addressing biomedical-related tasks, or introduce important biases in the data and the model due to the interpolation or the zero filling. Alternative methods to exploit the information of all the available subjects were proposed. A possible approach is the complementation of incomplete data representation, which consists in extracting a latent representation from both the complete and incomplete modalities, avoiding the need for data imputation [40, 41]. Most recent approaches aim at generating missing data either in the input space [42, 43, 44] or in an intermediate latent representation [45, 46, 47] relying on generative models such as Generative Adversarial Networks (GANs) [48, 49, 50, 42, 45, 46, 51] and their variants, or Variational Autoencoders [52]. This last approach has been successfully applied to the AD continuum investigation. It allows the exploitation of the availability of multimodal data for capturing the relationship among different data sources in the latent space, enabling the generation of one modality from another. Generative models, hence, appear as the route to be followed when dealing with missing data. Interestingly, the recently proposed Cycle-consistent GANs (Cycle-GANs) [53] have shown impressive results in various knowledge translation tasks since they offer a flexible and effective approach for learning mappings across different domains without relying on paired data.
@@
Moving to model interpretability, it is well established that with increasing model complexity, interpretability decreases drastically. However, in order to better understand the mechanisms that underlie the AD continuum and allow the models to be applied in clinical and real-life applications, it is vital to understand the reasons behind a certain output. In this case, eXplainable Artificial Intelligence (XAI), becomes essential to understand why a given model made a certain prediction. XAI encompasses post-hoc methods that allow the assignment of an importance score to each feature, reflecting its role in the classification task. This means finally opening the ‘black box’ of complex models hopefully increasing their exploitability in clinical practice. XAI is starting to be applied in multimodal frameworks to study neurodegenerative and psychiatric diseases. Few works could be found adopting simple gradient-based, feature perturbation methods, or attention weights [32, 49, 42, 54, 55, 56]. On top of this, another overlooked yet stringent aspect is the strong validation of XAI methods and the obtained explanations. The evaluation of explanation methods is still under-investigated, however, since explainability is meant to increase confidence in AI, it is vital to systematically analyze the obtained results referring also to the prior knowledge derived from the state-of-the-art.
@@
In this rapidly evolving landscape, where multimodality holds a central role, we aim at shading light on the importance of the complementary information that could be derived from advanced biomarkers such as brain FC and SNPs mutations, together with the well-established brain atrophy measures derived from sMRI, in detecting AD-related modulations. With this purpose, we will present a multimodal framework for AD detection and MCI conversion prediction, which addresses (i) heterogeneous data integration, (ii) missing data management, and (iii) interpretability. Regarding the former, our framework allows the integration of heterogeneous data featuring different dimensionality, e.g., 3D volumes and vectorized data, and nature, meaning from multi-domain, e.g., imaging and genetics. Concerning the second, we propose an approach that allows generating missing modalities in the latent space obtained after input feature reduction, ensuring high accuracy in reconstructing latent features while minimizing computational demands. The goal is to define a framework that can be generalized to any missing modality, without requiring to have at least one specific modality shared by all the subjects in the considered population. Finally, we aimed at emphasizing the interpretability of the proposed framework in order to reinforce its transparency and reliability by conducting an XAI post-hoc interpretability analysis, supplemented by a robust validation step, which enables to precisely discern the contribution of each input feature to the classification of subjects in the AD continuum, thus highlighting relevant phenotypic and genotypic biomarkers for AD.
2 Materials and Methods
@@ 2.1 Dataset
For each subject, the baseline timepoint was retained for all the selected imaging modalities, that is, sMRI and rs-fMRI. The respective genetic variants (SNPs) were then selected, if available. The final study cohort included $1911$ subjects, divided into healthy controls (CN), AD, MCI non-converter (MCInc), and MCI converter (MCIc). In detail, following the classification available on the ADNI website, an MCI subject was considered as MCIc if dementia was diagnosed at any timepoint after MCI diagnosis. Table 1 shows the Task 1 and Task 2 demographic information. For Task 1, the entire cohort, and the subset of individuals sharing both sMRI-fMRI and sMRI-SNPs are highlighted.
Table 1: Demographic information of the CN, AD, MCInc, and MCIc subjects. MCInc and MCIc have the following % of missing SNPs 59.6% and 30.8%, respectively, and the missing fMRI are 60.1% and 92%, respectively. MMSE = Mini-Mental State Examination; wild (W) and mutated (M) alleles in APOE4: 00 = homozygous (W/W), $1$ = heterozygous (W/M), $2$ = homozygous (M/M).
Task 1 Task 2 CN AD MCInc MCIc Cohorts All subjects sMRI-fMRI sMRI-SNPs All subjects sMRI-fMRI sMRI-SNPs Only for testing phase Count $644$ $321$ $253$ $332$ $66$ $152$ $646$ $289$ Age (y) $73.6\pm 6.6$ $72.4\pm 7.1$ $76.4\pm 5.4$ $75.1\pm 7.9$ $74.7\pm 8.1$ $75.8\pm 7.7$ $73\pm 8.5$ $74\pm 8.5$ Gender (F) 284 (360) 117 (204) 137 (116) 181 (151) 35 (31) 84 (68) 374 (272) 172 (117) MMSE $29.1\pm 1.1$ $29.1\pm 1.2$ $29.1\pm 1.1$ $23.1\pm 2.2$ $22.6\pm 2.6$ $23.3\pm 2.0$ $27.9\pm 1.8$ $26.9\pm 1.8$ APOE e4 ($0/1/2$) $455/168/21$ $224/87/10$ $189/57/7$ $109/156/67$ $19/34/13$ $49/76/27$ $377/205/64$ $103/143/43$
3D T1-w MRI and rs-fMRI acquisitions were considered as imaging input channels and were acquired with the following sequence parameters: (i) sagittal accelerated MPRAGE, TR/TE = shortest, TI = $900$ ms, flip angle = $9^{\circ}$, Field Of View = $256\times 256$ mm2, spatial resolution = $1\times 1\times 1$ mm3, slices = $176-211$), (ii) rs-fMRI: TR/TE =3000/303000303000/303000 / 30 ms, FA = $90^{\circ}$, FOV = $220\times 220\times 163$ mm3, $3.4$-mm isotropic voxel size. $200$ fMRI volumes were acquired in almost all subjects, with minimal variations in a small subset (e.g., $197$ or $195$ volumes). More details about the data acquisition can be found in [57]. Concerning the genetic data, DNA samples were genotyped using Illumina Human610-Quad or Illumina HumanOmniExpress BeadChip.